1. a process called “flash” electroplating, a current of 2.50 x 103 A passes through and electrolytic cell for 5.00 minutes. How many moles of electrons are driven through the cell?
1
Expert's answer
2020-11-02T09:02:39-0500

During 5.00 minutes ("t=" 5.00*60 = 300 seconds) a current "I" of 2.50x103 A passes through an electrolytic cell. This process transports the following charge "q" :

"q = It = 2.50\cdot10^3\text{ A}\cdot300\text{ s} = 7.5\cdot10^5" C.

According to the definition of the SI units, one coulomb corresponds exactly to 1/1.602176634x10-19 elementary charges:

"N = \frac{q}{1.602176634\cdot10^{-19}} = 4.7\cdot10^{24}" .

Finally, the number of the moles of electrons can be calculated as the ratio of the number of the electrons and the Avogadro constant:

"n(e^-) = \frac{N}{N_A} = \frac{4.7\cdot10^{24}}{6.02214076\cdot10^{23}\text{mol}^{-1}} = 7.77" mol.

Answer: 7.77 mol of electrons is driven through the cell during "flash" electroplating.